Suppose that four times the square of a number equals 20 times that number.What is the number?

Grade:11

3 Answers

Pawan Prajapati
3 years ago
Answer:

Answer is 5

explanation:

5*5=25

25*4=100

Therefore 20*5=100

Thank you !


Pawan Prajapati
Askiitian
Anshu Kumar
3 years ago
Suppose that number is x .Now we use the given condition then we get.
x^2*4 = 20*x
cancle one x from both side we get
4*x = 20
x = 20/4
x = 5
Ram Kushwah
3 years ago
Suppose the number is x
Four times the square of number=4x²
As per given 
4x²=20*x
4x²-20x=0
4x(x-5)=0
 
x=0 or x-5=0
Thus x=0 or 5
 
Neglecting the x=0 then x=5
 
Thus the said number=5